# MSP-MVS: Multi-Granularity Segmentation Prior Guided Multi-View Stereo

### v4

@@ Introduction
Multi-View Stereo (MVS) is a technique that employs images captured from different viewpoints to reconstruct 3D scenes or objects. Its diverse application spans across autonomous driving (Wei et al. 2023), augmented reality (Cao et al. 2020), and 3D printing (Zhu et al. 2021). Numerous ideas (Heise et al. 2013) and datasets (Yao et al. 2020) have significantly advanced its progress. However, it still faces the primary challenge of reconstructing large textureless areas.
@@
  • We achieve state-of-the-art performance on both the ETH3D dataset and Tanks & Temples dataset.
@@
MVSNet (Yao et al. 2018) is the first to utilize differentiable 3D cost volumes to construct deep neural networks, thereby becoming the foundation for the abundant subsequent methods. To reduce memory, Cas-MVSNet (Yang et al. 2020) adopts a coarse-to-fine strategy to retain depths across multiple scales. IterMVS-LS (Wang et al. 2022a) leverages a lightweight probability estimator to encode depth distribution during regularization. For feature extraction, AAR-MVSNet (Wei et al. 2021) employed deformable convolutions to achieve adaptive feature aggregation. Differently, MVSTER (Wang et al. 2022b) leverages the transformer structure to introduce the multi-head attention mechanism. Concerning geometrical topology, Geo-MVSNet (Zhang et al. 2023b) adopts a two-branch geometry fusion network to enhance geometry perception. Regarding surface topology, RA-MVSNet (Zhang et al. 2023a) associates hypothesis planes with surface patches to enhance perception field. Moreover, Raymvsnet++ (Shi et al. 2023) applies an epipolar transformer to enhance feature aggregation along camera rays. GoMVS (Wu et al. 2024) integrates a geometrically consistent propagation module to refine cost aggregation. Differently, our method targets on aggregating multi-granularity depth edges for edge-confined patch deformation. Despite this, most of them still suffer from either unaffordable memory costs or limited generalization.
@@
Table 3: Quantitative ablation studies on ETH3D dataset.

### v5

@@ Introduction
Multi-View Stereo (MVS) is a technique that employs images captured from different viewpoints to reconstruct 3D scenes or objects. Its diverse application and downstream task spans across object detection (Lu et al. 2023, 2024), image denoising (Yang et al. 2024; Li et al. 2022), video representation (Liu et al. 2024a, b), multiview clustering (Guan et al. 2024b, a), etc. Numerous ideas (Chen et al. 2024a, b; Kang et al. 2025) and datasets (Yao et al. 2020) have advanced its progress. However, it still faces the primary challenge of reconstructing large textureless areas.
@@
Recently, the advancement of deep learning has led to the emergence of numerous large-language models (Luo et al. 2024a, b), gaussian splatting algorithms (Gan et al. 2023a, b), and self-supervised learning methods (Li and Zhang 2022; Song et al. 2025; Chen et al. 2023, 2024c; Cheng and Sun 2024; Sun et al. ) and vision-language model (Chen et al. 2024d; Qian et al. 2024; Chen et al. 2024e). In MVS field, MVSNet (Yao et al. 2018) is the first to utilize differentiable 3D cost volumes to construct deep neural networks. To reduce memory, Cas-MVSNet (Yang et al. 2020) adopts a coarse-to-fine strategy to retain depths across multiple scales. IterMVS-LS (Wang et al. 2022a) leverages a lightweight probability estimator to encode depth distribution during regularization. For feature extraction, AAR-MVSNet (Wei et al. 2021) employed deformable convolutions to achieve adaptive feature aggregation. Differently, MVSTER (Wang et al. 2022b) leverages the transformer structure to introduce the multi-head attention mechanism. Geo-MVSNet (Zhang et al. 2023b) adopts a two-branch geometry fusion network to enhance geometry perception. RA-MVSNet (Zhang et al. 2023a) associates hypothesis planes with surface patches to enhance perception field. Moreover, Raymvsnet++ (Shi et al. 2023) applies an epipolar transformer to enhance feature aggregation along camera rays. GoMVS (Wu et al. 2024) integrates a geometrically consistent propagation module to refine cost aggregation. Differently, our method targets on aggregating multi-granularity depth edges for edge-confined patch deformation. Despite this, most of them still suffer from either unaffordable memory costs or limited generalization.
@@
Table 3: Quantitative results of the ablation studies on ETH3D benchmark to validate each proposed component.